I need help learning the steps to this problem.
For the following polynomial find f(t-2)
((( f(x)=-x^2+4x=6 )))

Answer by stanbon(75887)

find f(t-2) if f(x)=-x^2+4x+6
---------
f(t-2)=(t-2)^2+4(t-2)+6
f(t-2)=t^2-4t+4+4t-8+6
f(t-2)=t^2+2
Cheers,
Stan H.